### v1

# Beijing 100084, China

Abstract
In this paper, we propose a novel function named Rate Distortion-in-Distortion (RDD) function as an extension of the classical rate-distortion (RD) function, where the expected distortion constraint is replaced by the Gromov-type distortion. This distortion, integral to the Gromov-Wasserstein (GW) distance, effectively defines the similarity in spaces of different dimensions without a direct metric between them. While our RDD function qualifies as an informational RD function, encoding theorems substantiate its status as an operational RD function, thereby underscoring its potential applicability in real-world source coding. Due to the high computational complexity associated with Gromov-type distortion, the RDD function cannot be solved analytically. Consequently, we develop an alternating mirror descent algorithm that significantly reduces computational complexity by employing decomposition, linearization, and relaxation techniques. Simulations on classical sources and different grids demonstrate the effectiveness of our algorithm. By examining the distinctions and connections between the RDD function and the RD function, we anticipate that RDD function will play a novel role in foreseeable future scenarios.
Index Terms:
Rate-distortion, Gromov-Wasserstein metric.
I Introduction
In this paper, we propose a new framework named Rate Distortion-in-Distortion (RDD) function, which originates from the Rate-Distortion (RD) [1, 2] theory and incorporates concepts from the Gromov-Wasserstein (GW) [3, 4] distance. Specifically, the RDD function adopts mutual information as the objective function, similar to the RD function, while replacing the expected distortion constraint with the Gromov-type distortion [5] of two measurable spaces as used in the GW distance [6].
The RDD function serves as an extension of the classical RD function. Like other RD extensions, it incorporates structural considerations into the rate-distortion problem, such as the perception term introduced by RDP function in the field of lossy compression [7]. Specifically, RDD adopts the Gromov-type distortion of two measurable spaces, rather than solely quantifying distances of the same dimension. This insight is provided by the GW distance, which offers a flexible and comprehensive framework for evaluating information processing systems, addressing limitations [8] in RD theory: the difficulty in defining a distortion measure across spaces of different dimensions, and the challenge of dealing with unknown distortion measures even within spaces of the same dimension.
Our exploration of GW distance is motivated by the similarities between the Optimal Transport (OT) problem [9] and the RD problem. Specifically, the RD objective function partially corresponds to the entropy regularization in regularized OT, and the constraints on conditional probability in the RD function are related to the marginal distributions in OT problem [10]. As an important model in OT theory, GW distance provides a powerful method for evaluating discrepancies across distributions by measuring differences between distances rather than individual points [3]. Notably, the GW distance offers significant flexibility by enabling the comparison of probability distributions. In scenarios where the underlying metric spaces differ, GW distance captures structural and relational information that traditional distances may overlook [6]. Furthermore, when the distributions are supported on identical metric spaces, GW distance captures intrinsic geometric properties, thus offering valuable additional insights, such as shape analysis. Such functionality makes GW distance applicable in a wide range of mathematical and computational areas, including graph matching [6, 11, 12, 13], natural language processing [14], and machine learning [15, 16]. Although the RDD function, integrated with Gromov-type distortion, currently lacks specific application scenarios, we anticipate it would play a relevant role in the field of information theory.
The RDD function is defined with mutual information as the objective function, positioning it naturally as an informational rate-distortion function. By invoking the coding theorem presented in [17], we can show that the RDD function corresponds to the minimum rate required for a specific coding task. Consequently, the RDD function holds operational significance for assessing the performance of real-world source coding.
It is noteworthy that the RDD function involves quadratic constraints from the Gromov-type distortion, and thus constructing analytical solutions becomes vastly challenging. Therefore, an effective numerical algorithm is in high demand. In order to effectively reduce the quadratic arithmetic complexity, we develop an alternating mirror descent numerical algorithm, drawing inspiration from the computation of GW distance to decompose the Gromov-type distortion [18]. In addition, the RDD function involves an unknown marginal distribution, necessitating the relaxation of certain constraints and linearization of conditional probabilities to avoid complicated iterative operations caused by dual parameters in the Lagrangian function. By adapting the aforementioned techniques, we are able to reduce the arithmetic complexity from quartic to cubic.
Moreover, we examine the proposed numerical algorithm on classical sources, within the Euclidean metric spaces, including Gaussian, uniform, and Laplacian sources. Furthermore, the algorithm has also been evaluated using both uniform and non-uniform grid points across various dimensions. Finally, the distinctions and connections between RDD and RD functions are examined by the properties of a function that incorporates both Gromov-type distortion and classical distortion.
II Rate Distortion-in-Distortion Function
Let $(\mathcal{X},d_{\mathcal{X}},P_{\mathcal{X}})$ denote a metric measure space, and let $(\mathcal{Y},d_{\mathcal{Y}})$ represent a metric space, where $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}^{+}$ and $d_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ are two distances, while $P_{\mathcal{X}}$ is a probability measure on $\mathcal{X}$. We use $X$ and $Y$ to represent the random variables of spaces $\mathcal{X},\mathcal{Y}$, respectively. Clearly, the probability distribution of $X$ is $P_{\mathcal{X}}$, and for convenience, we will use $P_{X}$ to represent the distribution of $X$ throughout this paper. Given the unknown probability distribution on space $\mathcal{Y}$, the joint distribution of random variables $X$ and $Y$ refers to the product of the marginal distribution $P_{X}$ and the conditional probability $P_{Y|X}$, i.e., $P_{XY}=P_{X}\cdot P_{Y|X}$. We hereby introduce the RDD function as:
$R_{G}(D)=\min\limits_{P_{Y|X}}\quad I(X;Y)$ (1a)
s.t. $\mathbb{E}\left[f(d^{q}_{\mathcal{X}}(X,X^{\prime}),d^{q}_{\mathcal{Y}}(Y,Y^{\prime}))\right]\leq D,$
$\quad(X,Y,X^{\prime},Y^{\prime})\sim P_{XY}\times P_{XY},$ (1b)
where $q\geq 1$, and $I(X;Y)$ denotes the mutual information between $X$ and $Y$. The function $f$ measures the difference between the two metric $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$. Meanwhile, the joint distribution of $(X,Y,X^{\prime},Y^{\prime})$ is regarded as the product of two independent and identical distributions. Consequently, the expectation $\mathbb{E}$ in (1b) is taken with respect to the product of the joint distribution $P_{XY}\times P_{XY}$.
The commonly used element-wise loss function [13] $f$ includes Euclidean distance (MSE) $f(a,b)=(a-b)^{2}$ and KL-divergence $f(a,b)=a\log\frac{a}{b}-a+b$. In the rest of the paper, we adopt Euclidean distance to discuss theoretical properties and numerical algorithms due to its popularity in information theory. In such a case, the RDD function can be written as:
$R_{G}(D)=\min\limits_{P_{Y|X}}\quad I(X;Y)$ (2a)
s.t. $\mathbb{E}\left[|d^{q}_{\mathcal{X}}(X,X^{\prime})-d^{q}_{\mathcal{Y}}(Y,Y^{\prime})|^{2}\right]\leq D,$
$\quad(X,Y,X^{\prime},Y^{\prime})\sim P_{XY}\times P_{XY}.$ (2b)
Notably, the constraint in (2b) now becomes the Gromov-type distortion $\mathcal{E}(d^{q}_{\mathcal{X}},d^{q}_{\mathcal{Y}},P_{Y|X})$,where $P_{XY}=P_{X}\cdot P_{Y|X}$ [5]:
$\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}\!\!\!|d_{\mathcal{X}}^{q}(x,x^{\prime})-d_{\mathcal{Y}}^{q}(y,y^{\prime})|^{2}\mathop{}\!\mathrm{d}P_{XY}(x,y)\mathop{}\!\mathrm{d}P_{XY}(x^{\prime},y^{\prime}),$
Here we notice the similarity between the RDD function and the classical RD function. The classical RD function is obtained by minimizing the mutual information between the source and the reproduction subject to an average distortion constraint. Concretely, given a source random variable $X\in\mathcal{X}$ with probability distribution $P_{X}$, a reproduction $Y\in\mathcal{Y}$, the RD function is defined as [19, 1]:
$R(D)=\min\limits_{P_{Y|X}:\ \mathbb{E}[d(X,Y)]\leq D}{I(X;Y)}.$ (3)
The distortion measure $d:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ here is primarily effective for spaces of identical dimensions. We refer to our function as rate distortion-in-distortion because we replace the distortion measure $d$ in RD function by the distortion-in-distortion $(d^{q}_{\mathcal{X}}-d^{q}_{\mathcal{Y}})^{2}$, which is inspired by the GW distance [3]:
$\mathrm{GW}^{2}_{q}((X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}}))=$
$\inf_{P_{XY}}\int_{(X\times Y)^{2}}\!\!\!\!\!\!\!\!\!\!\!\!(d^{q}_{\mathcal{X}}(x,x^{\prime})-d^{q}_{\mathcal{Y}}(y,y^{\prime}))^{2}\,\mathrm{d}P_{XY}(x,y)\,\mathrm{d}P_{XY}(x^{\prime},y^{\prime})$
The GW distance concerns the distance between two metric measure spaces, while the RD function focuses on compression. Leveraging insights from GW distance, the RDD function extends the application and definition of the RD problem, being able to handle spaces of different dimensions.
Similar to those in RD theory, we present a simple yet useful property of the proposed RDD function.
Proposition 1.
RDD function $R_{G}(D)$ is zero if and only if
$D\geq D_{max}:=\min\limits_{P_{Y}}\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}|d^{q}_{\mathcal{X}}(x,x^{\prime})-d^{q}_{\mathcal{Y}}(y,y^{\prime})|^{2}\cdot\\ P_{X}(x)P_{X}(x^{\prime})P_{Y}(y)P_{Y}(y^{\prime})\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}x^{\prime}\mathop{}\!\mathrm{d}y^{\prime}.$ (4)
Proof:
See Appendix due to the space limitation. ∎
III Coding Theorem
In this section, we use a coding theorem to show that $R_{G}(D)$ is the minimum number of bits required by a coding task for each source symbol. To start with, we define a code sequence as $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$, where encoding function $f_{n}:\mathcal{X}^{n}\times\mathbb{R}\rightarrow\mathbb{Z}^{+}$ and decoding function $\varphi_{n}:\mathbb{Z}^{+}\times\mathbb{R}\rightarrow\mathcal{Y}^{n}$. Each $f_{n}$ represents an encoder, and each $\varphi_{n}$ represents a decoder. For two random vectors $(X_{1},\cdots,X_{n})$ and $(Y_{1},\cdots,Y_{n})$, they are said to satisfy the Gromov-type distortion constraint regarding distortion $D$, if
$\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!|d^{q}_{\mathcal{X}}(x,x^{\prime})-d^{q}_{\mathcal{Y}}(y,y^{\prime})|^{2}\mathop{}\!\mathrm{d}P_{X_{i}Y_{i}}^{i}(x,y)\mathop{}\!\mathrm{d}P_{X_{i}Y_{i}}^{i}(x^{\prime},y^{\prime})\leq D$
satisfies for all positive integers $i\leq n$, where $P_{X_{i},Y_{i}}^{i}$ represents the joint distribution on the $i$-th variable pair $(X_{i},Y_{i})$. Now let $X_{1}$, $X_{2}$, $\cdots$, $X_{n}$ be independent and $P_{X}$-distributed random variables. Using the Theorem 3 in [17], we have the following coding theorem.
Theorem 1.
For real numbers $R$ and $D$, $R\geq R_{G}(D)$ is a sufficient and necessary condition for the existence of a code sequence $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$ and a sequence $\{U_{n}\}_{n=1}^{\infty}$ of random variables such that
$\limsup_{n\to\infty}\frac{1}{n}H(f_{n}(X_{1},\cdots,X_{n},U_{n})|U_{n})\leq R,$ (5)
where $U_{n}$ and $\{X_{i}\}_{i=1}^{\infty}$ are independent for all $n\in\mathbb{Z}^{+}$, and $(X_{1},\cdots,X_{n})$ and $\varphi_{n}(f_{n}(X_{1},\cdots,X_{n},U_{n}),U_{n})$ satisfy the Gromov-type distortion constraint regarding $D$ for all $n\in\mathbb{Z}^{+}$.
Proof:
See Appendix due to the space limitation. ∎
The random variables $U_{1}$, $U_{2}$, $\cdots$, $U_{n}$ in the coding theorem serve as common randomness between the encoders and decoders. Since we can represent $f_{n}(X_{1},\cdots,X_{n},U_{n})$ using a variable-length code whose average length is approximately $H(f_{n}(X_{1},\cdots,X_{n},U_{n})|U_{n})$, we need at least $R_{G}(D)$ bits per source symbol to ensure that the Gromov-type distortion constraint regarding $D$ is satisfied.
The coding theorem presented in this section is more suitable for practical encoding design scenarios. Additionally, we provide an alternative perspective on the coding theorem in the Appendix.
IV Alternating Mirror Descent Algorithm
Notably, the quadratic nature of the Gromov-type distortion constraint makes obtaining a closed-form solution for the RDD function theoretically challenging, even for Gaussian sources. To compute the RDD function in practical applications, effective algorithms are essential. However, the quartic arithmetic complexity from Gromov-type distortion presents significant computational challenges. To address these difficulties, we adapt the widely used Lagrangian multipliers method [10] with certain effective techniques to develop a novel numerical algorithm for the computation of the RDD function, which is referred to as Alternating Mirror Descent (AMD) algorithm.
In the sequel, we adopt a discretized formulation of the RDD function with discrete sources for computational convenience. Consider a memoryless source $X\in\mathcal{X}$ with reproduction $Y\in\mathcal{Y}$, where $\mathcal{X}\in\left\{x_{1},\cdots,x_{M}\right\},\mathcal{Y}\in\left\{y_{1},\cdots,y_{N}\right\}$ are finite and discrete alphabets. We denote $w_{ij}=P_{Y|X}(x_{i},y_{j})$ and $W$ is the matrix whose $(i,j)$ element is $w_{ij}$. Denote $D^{\mathcal{X}},D^{\mathcal{Y}}$ as the matrix whose $(i,j)$ element is $d_{\mathcal{X}}^{q}(x_{i},x^{\prime}_{j}),\ d_{\mathcal{Y}}^{q}(y_{i},y^{\prime}_{j})$ respectively. Let $p_{i}=P_{X}(x_{i}),r_{j}=P_{Y}(y_{j})$ and denote $E_{ij}:=D^{\mathcal{X}}_{ij}p_{j},C_{ij}:=E_{ij}p_{i}$ as the metrics for short. By introducing an auxiliary variable $\bm{r}$, the RDD function can be written in the discrete form as:
$\min\limits_{w_{ij}\geq 0,r_{j}\geq 0}\;\sum\limits_{i=1}^{M}{\sum\limits_{j=1}^{N}{(w_{ij}p_{i})\left[\ln w_{ij}-\ln r_{j}\right]}}$ (6a)
[…]
$\quad\quad\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)\leq D,$ (6c)
where the Gromov-type distortion $\mathcal{E}$ can be written as:
$\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)=\sum\limits_{i,i^{\prime},j,j^{\prime}}|D^{\mathcal{X}}_{ii^{\prime}}-D^{\mathcal{Y}}_{jj^{\prime}}|^{2}w_{ij}w_{i^{\prime}j^{\prime}}p_{i}p_{i^{\prime}}.$
IV-A Equivalent Transformation
Intuitively, the cost of evaluating (6c) is often unacceptable in the field of information science. Following the approach
in [12, 18], we can decompose the function $\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)$ into one constant term $\widetilde{\mathcal{C}}_{1}$ and two quadratic terms of $W$ [^1] Unlike the GW distance, in the RDD function, one of the marginal distributions is unknown, leading to the decomposition of the Gromov-type distortion containing only a single constant term.:
$\widetilde{\mathcal{C}}_{1}=P_{X}^{\top}(D^{\mathcal{X}})^{\odot 2}P_{X},\;\widetilde{\mathcal{C}}_{2}=(W^{\top}P_{X})^{\top}(D^{\mathcal{Y}})^{\odot 2}W^{\top}P_{X},$
$\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)=\widetilde{\mathcal{C}}_{1}+\widetilde{\mathcal{C}}_{2}-2\left\langle CWD^{\mathcal{Y}},W\right\rangle,$
[…]
Moreover, similar to the treatment in [20], the constraints on marginal distributions in the second term of (6c), expressed as $\sum_{i=1}^{M}w_{ij}p_{i}=r_{j},\forall j$, can be relaxed during the subsequent Lagrangian analysis. This relaxation not only simplifies the formulation of the problem during the Lagrangian analysis, but also leads to a closed-form solution in the $r_{j}$ direction.
Following this way, we propose a semi-relaxed RDD function, i.e.,
$\min\limits_{w_{ij}\geq 0,r_{j}\geq 0}\;$ $f(W,\bm{r})\triangleq\sum\limits_{i=1}^{M}{\sum\limits_{j=1}^{N}{(w_{ij}p_{i})\left[\ln w_{ij}-\ln r_{j}\right]}}$ (7a)
s.t. $\sum\limits_{j=1}^{N}{w_{ij}}=1,\;\sum\limits_{j=1}^{N}{r_{j}}=1,\;\forall i,$ (7b)
$\widetilde{\mathcal{C}}_{1}+\widetilde{\mathcal{C}}_{2}-2\left\langle CWD^{\mathcal{Y}},W\right\rangle\leq D.$ (7c)
Theorem 2.
The optimal solution to the semi-relaxed RDD function (7) is exactly that to the original RDD function (6).
Proof:
See Appendix due to the space limitation. ∎
IV-B Algorithm Derivation and Implementation
By introducing the multipliers $\bm{\alpha}\in\mathbb{R}^{M},\lambda\in\mathbb{R}^{+},\eta\in\mathbb{R}$, the Lagrangian of the semi-relaxed RDD function (7) is:
$\mathcal{L}(W,\bm{r};\bm{\alpha},\lambda,\eta)=\sum\limits_{i=1}^{M}{\sum\limits_{j=1}^{N}{(w_{ij}p_{i})\left[\ln w_{ij}-\ln r_{j}\right]}}$
$+\sum\limits_{i=1}^{M}{\alpha_{i}\!\!\left(\sum\limits_{j=1}^{N}{w_{ij}\!-\!1}\right)}\!+\!\eta\!\left(\sum\limits_{j=1}^{N}{r_{j}}\!-\!1\right)-2\lambda\left\langle CWD^{\mathcal{Y}},W\right\rangle$
$+\lambda\left(P_{X}^{\top}(D^{\mathcal{X}})^{\odot 2}P_{X}+(W^{\top}P_{X})^{\top}(D^{\mathcal{Y}})^{\odot 2}W^{\top}P_{X}\right)-\lambda D.$
Our key ingredient is optimizing the primal variables $W,\bm{r}$ in an alternative manner. Based on the Lagrangian of the RDD function, we take derivatives with respect to the primal variables $W$ and $\bm{r}$ for its optimal expression.
a) Updating $W=(w_{ij})$: When taking the derivative of $\mathcal{L}(W,\bm{r};\bm{\alpha},\lambda,\eta)$ with respect to the primal variable $W$, to address the first-order condition, we utilize the mirror descent method to linearize the quadratic component of $W$, one has
$\frac{\partial\mathcal{L}}{\partial w_{ij}}=p_{i}\left[1+\ln w_{ij}-\ln r_{j}\right]+\alpha_{i}$
$-4\lambda\sum\limits_{i^{\prime}=1}^{M}{\sum\limits_{j^{\prime}=1}^{N}}{C_{ii^{\prime}}D^{\mathcal{Y}}_{jj^{\prime}}w^{(k)}_{i^{\prime}j^{\prime}}}+2\lambda\sum\limits_{i^{\prime}=1}^{M}{\sum\limits_{j^{\prime}=1}^{N}}{(D_{j,j^{\prime}}^{\mathcal{Y}})^{2}p_{i}p_{i^{\prime}}w^{(k)}_{i^{\prime}j^{\prime}}}.$
Thus $w_{ij}^{(k+1)}$ can be expressed in the closed-form of
$w_{ij}^{(k+1)}=\frac{r_{j}^{(k)}\exp{\big{(}\lambda\sum\limits_{i^{\prime},j^{\prime}}{w^{(k)}_{i^{\prime}j^{\prime}}(4E_{ii^{\prime}}D^{\mathcal{Y}}_{jj^{\prime}}-2(D_{j,j^{\prime}}^{\mathcal{Y}})^{2}p_{i^{\prime}})}\big{)}}}{\sum\limits_{l=1}^{N}{r_{l}^{(k)}\exp{\big{(}\lambda\sum\limits_{i^{\prime},j^{\prime}}{{w^{(k)}_{i^{\prime}j^{\prime}}(4E_{ii^{\prime}}D^{\mathcal{Y}}_{lj^{\prime}}-2(D_{l,j^{\prime}}^{\mathcal{Y}})^{2}p_{i^{\prime}})}\big{)}}}}}.$
b) Updating $\bm{r}$: Taking the derivative of $\mathcal{L}(W,\bm{r};\bm{\alpha},\lambda,\eta)$ with respect to the primal variable $\bm{r}$ leads to the following equation
$\frac{\partial\mathcal{L}}{\partial r_{j}}=-\sum\limits_{i=1}^{M}{w_{ij}p_{i}\frac{1}{r_{j}}+\eta},$
which implies $r_{j}=\left(\sum_{i=1}^{M}w_{ij}p_{i}\right)/\eta$. Substituting the obtained expression into the constraint of $\bm{r}$, we have $\sum_{j=1}^{N}{\left[\sum_{i=1}^{M}w_{ij}p_{i}/\eta\right]}=1.$ Since $\sum_{i=1}^{M}\sum_{j=1}^{N}p_{i}w_{ij}=1$, one obtains $\eta=1$. Then we can update $\bm{r}$ by $r_{j}=\sum\limits_{i=1}^{M}{w_{ij}p_{i}}.$
To summarize, we update the variables $W$ and $\bm{r}$ in an alternating manner. For clarity, the pseudo-code is presented in the below Algorithm 1.
Algorithm 1 Alternating Mirror Descent (AMD) algorithm
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric space $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrix $D^{\mathcal{X}},D^{\mathcal{Y}}$
1:  Initialization: $\lambda,r_{j}=1/N$
2:  Set $E_{ij}:=D^{\mathcal{X}}_{ij}p_{j}$, $W^{(0)}=(\bm{1}_{M}\bm{1}_{N}^{\top})/N$
3:  for $k=0:max\_iter$ do
4:     Let $F_{ij}=\exp{\big{(}4\lambda\sum_{i^{\prime},j^{\prime}}{w^{(k)}_{i^{\prime}j^{\prime}}E_{ii^{\prime}}D^{\mathcal{Y}}_{jj^{\prime}}\big{)}}}$
5:     Let $H_{ij}=\exp{\big{(}-2\lambda\sum_{i^{\prime},j^{\prime}}{w^{(k)}_{i^{\prime}j^{\prime}}(D_{j,j^{\prime}}^{\mathcal{Y}})^{2}p_{i^{\prime}}\big{)}}}$
6:     Update $w^{(k+1)}_{ij}\leftarrow r_{j}^{(k)}F_{ij}H_{ij}/\sum_{l=1}^{N}{r_{l}^{(k)}F_{il}H_{il}}$
7:     Update $r_{j}^{(k+1)}\leftarrow\sum_{i=1}^{M}w_{ij}^{(k+1)}p_{i}$
7:  $\sum_{i,j}(w_{ij}p_{i})\left[\ln w_{ij}-\ln r_{j}\right]$
The computational complexity of the proposed algorithm is $\mathcal{O}(M^{2}N+MN^{2}))$. The primary bottleneck occurs in line 5, which involves matrix multiplication. In contrast, all other operations consist solely of vector multiplications.
V Numerical Experiments and Discussions
This section investigates the relevant properties of the RDD function through numerical experiments. For starters, we compute the RDD function of the discrete versions of three classical distributions (i.e., the Gaussian source, Laplacian source and uniform source) across various metric spaces. All three sources utilize the square of $L_{2}$ norm distance, i.e., $q=2$. For simplicity, we refer to $dim(\mathcal{X})$ as the dimension of the metric space $\mathcal{X}$. For the implementation of the proposed AMD algorithm, the multiplier $\lambda$ is chosen as an arithmetic sequence of $100$ elements and we compute the result for each given $\lambda$. The parameter $max\_iter$ is set to $100$.
First, we consider the case where points in the metric measure spaces $\mathcal{X}$ and $\mathcal{Y}$ are distributed in a uniform grid. Specifically, we truncate the sources of each dimension into an interval $[-h,h]$ and then discretize the interval by a set of uniform grid points: $x_{i}=-h+\delta/2+(i-1)\cdot\delta$, $\delta=2\cdot h/K$, $i=1,\cdots,K,$ where $K$ denotes the number of coordinates of each dimension. Therefore, the two-dimensional and three-dimensional metric spaces contain $K^{2}$ and $K^{3}$ points, respectively. The probability distribution $P_{X}$ is generated from the density functions of sources and then normalized. In the sequel, we take $h=8$ for all three sources.
Figure 1: The RDD functions of different sources and different metric space dimensions. Left: $dim(\mathcal{X})=dim(\mathcal{Y})=1$; Right: $dim(\mathcal{X})=2,dim(\mathcal{Y})=3$.
It can be observed that the proposed AMD algorithm demonstrates the convergence property to some extent for discrete sources.
Next, we simulate the point distributions on spherical surfaces i.e., a circle and a sphere, each with a radius of $4$ units. For the circle, we calculate the angular increments to evenly distribute $20$ points around its perimeter and use trigonometric functions to obtain their Cartesian coordinates. As for the sphere, we establish the angular positions for $20$ points across each latitude and longitude, employing a meshgrid to create a grid that covers the sphere’s surface comprehensively. This approach yields a total of $20$ points on a circle and $400$ points on a sphere, respectively. The corresponding RDD curve is plotted in Fig. 2. This illustration verifies that the proposed RDD function and AMD algorithm extend their applicability beyond merely uniform grids.
Finally, we discuss the relationship between RDD and RD functions by considering the following problem, i.e., the constraints involve both distortion and distortion-in-distortion,
$R(D;\theta)=\min\limits_{P_{Y|X}}\quad I(X;Y)$ (8a)
s.t. $\theta\mathbb{E}\left[|d^{q}_{\mathcal{X}}(X,X^{\prime})-d^{q}_{\mathcal{Y}}(Y,Y^{\prime})|^{2}\right]$
$+(1-\theta)\mathbb{E}[d(X,Y)]\leq D,(X,Y,X^{\prime},Y^{\prime})\sim P_{XY}\times P_{XY},$ (8b)
Figure 2: The RDD function of Gaussian source on spherical surfaces. The dimensions of metric spaces are $dim(\mathcal{X})=2,dim(\mathcal{Y})=3$.
The constraint (8b) draws inspiration from the form of Fused GW [21], which combines Wasserstein and Gromov-Wasserstein distances to jointly account for the features and structures of graphs. The function (8) can be solved by slight modifications to the proposed AMD method introduced in Section IV while incorporating the idea of the Blahut-Arimoto (BA) algorithm [22]. The detailed algorithmic procedure is provided in the Appendix . Here, we consider a one-dimensional Gaussian distribution with $\sigma=2$ and set the distortion measure $d$ as the squared error.
Figure 3: The $R(D;\theta)$ surface and its corresponding cross-section curves. Left: $R(D;\theta)$ surface with $\theta\in[0,0.1]$; Middle: $R$-$D$ curve; Right: $R$-$\theta$ curve. The dimensions of metric spaces are $dim(\mathcal{X})=dim(\mathcal{Y})=1$.
Fig. 3 illustrates the $R(D;\theta)$ surface along with its corresponding cross-section curves. Whether we examine the overall surface (left), the $R$-$D$ curve with a fixed $\theta$ (middle), or the $R$-$\theta$ curve with a fixed D (right), it is apparent that as $\theta$ increases (indicating a growing influence of distortion-in-distortion) and $D$ decreases, the $R$-value increases monotonically. This indicates that the impact of the distortion-in-distortion term on the distortion term is significant. We acknowledge that the distortion-in-distortion term lacks theoretical convexity. However, even when $\theta$ is very small (middle), where the convex distortion term dominates, the results of numerical experiments clearly demonstrate the monotonic relationship of $R$ concerning both $\theta$ and $D$. Due to space limitations, the conditional probabilities $P_{Y|X}$ for various values of $\theta$ and $D$ are presented in Appendix . Additionally, we can observe how changes in $\theta$—specifically, changes in the weight of the distortion-in-distortion term—affect the conditional probability $P_{Y|X}$.
VI Conclusion
This paper introduces a novel function i.e., RDD function, which serves as an extension of the classical RD function. The theoretical validity of the RDD function is established through a coding theorem, demonstrating its rationale as an operational rate-distortion function. In order to conquer the high computational complexity caused by the Gromov-type distortion and to efficiently compute the RDD function, we propose a low-complexity AMD algorithm and test it on several classical sources for various metric spaces. This work primarily presents a theoretical exploration, while its practical applications remain under investigation. As an extension of the classical RD function—just like the RDP function—we believe that the RDD function also holds promise for applications in a wide range of fields, such as the point cloud compression.
Supplement to Fig. 3
Here, we present the conditional probability $P_{Y|X}$ for different values of $\theta$ and $D$ in Fig. 5, and select several fixed values of $X$ to provide a more intuitive demonstration.
Figure 5: The conditional probability $P_{Y|X}$ for different $\theta$ and $D$. Left: $D=1$; Right: $D=3$; Upper: $\theta=0$; Middle: $\theta=0.01$; Lower: $\theta=0.02$.
For $\theta=0$, the corresponding distribution represents the solution to the classical RD problem. It can be observed that, as $\theta$ increases, the conditional distribution progressively deviates from that of the classical RD solution, which indicates the significant impact of the distortion-in-distortion term on the distortion term.
Algorithm 2 AMD algorithm for the computation of problem (8)
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric space $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrix $D^{\mathcal{X}},D^{\mathcal{Y}}$
1:  Initialization: $\lambda,\theta,r_{j}=1/N$
2:  Set $E_{ij}:=D^{\mathcal{X}}_{ij}p_{j}$, $W^{(0)}=(\bm{1}_{M}\bm{1}_{N}^{\top})/N$
3:  for $k=0:max\_iter$ do
4:     Let $F_{ij}=\exp{\big{(}4\lambda\theta\sum_{i^{\prime},j^{\prime}}w^{(k)}_{i^{\prime}j^{\prime}}E_{ii^{\prime}}D^{\mathcal{Y}}_{jj^{\prime}}\big{)}}$
5:     Let $H_{ij}=\exp{\big{(}-2\theta\lambda\sum_{i^{\prime},j^{\prime}}w^{(k)}_{i^{\prime}j^{\prime}}(D_{j,j^{\prime}}^{\mathcal{Y}})^{2}p_{i^{\prime}}\big{)}}$
6:     Let $G_{ij}=\exp(-\lambda(1-\theta)d_{ij})$
7:     Update $w^{(k+1)}_{ij}\leftarrow r_{j}^{(k)}F_{ij}H_{ij}G_{ij}/\sum_{l=1}^{N}{r_{l}^{(k)}F_{il}H_{il}G_{il}}$
8:     Update $r_{j}^{(k+1)}\leftarrow\sum_{i=1}^{M}w_{ij}^{(k+1)}p_{i}$
8:  $\sum_{i,j}(w_{ij}p_{i})\left[\ln w_{ij}-\ln r_{j}\right]$